Find the derrivative bellow
dz (a) Suppose z = f(x,y) = xexy. Compute if x(t) if x(t) = f and y(t) = +-1.
Answer #1

Given z = f(x1y) ay alt) = 2 and y) = t - dz dt f da dt + 20 ty dy - of and fy - by of re he (Gas) e ree te te? ta (geply) Redz f da dt + fy dy dt X(t) = 42 and 9 Lt) - et x = t² and y = dz dt le collis) + f (012) dz. (et + 2 ) Cat) & (6) et (2/2) dt